OpenStudy (anonymous):

factor the expression x^3+216

OpenStudy (anonymous):

this is a sum of cubes, recall: a^3+b^3=(a+b)(a^2-ab+b^2)

OpenStudy (anonymous):

(x^(3)+216) Since both terms are perfect cubes, the binomial can be factored using the sum of cubes formula: a^(3)+b^(3)=(a+b)(a^(2)-ab+b^(2)). (x+6) ( x^2-6x+36)